Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 35,000 helmets, using 22,500 kilograms of plastic. The plastic cost the company $171,000.

     According to the standard cost card, each helmet should require 0.6 kilograms of plastic, at a cost of $8 per kilogram.

Required:
1.

According to the standards, what cost for plastic should have been incurred to make 35,000 helmets? How much greater or less is this than the cost that was incurred? (Round Standard kilograms of plastic per helmet to 1 decimal place.)

     

2.

Break down the difference computed in (1) above into a materials price variance and a materials quantity variance. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)
Answer #1

1)

Number of helmets 35000
Standard quantity per helmet 0.60
Standard kg allowed 21000
Standard cost per kg 8
Total Standard cost 168000
Actual Cost of material 171000
Total Standard cost of material 168000
Total material variance 3000 U

2) Material price variance = (8*22500-171000) = 9000 F

Material quantity variance = (21000-22500)*8 = 12000 U